# Structured Probabilistic Inference & Generative Modeling

### v1

@@ 1 Introduction
When a terrorist attack struck the Boston Marathon in 2013, emergency responders faced around 250 casualties within minutes, overwhelming the city’s trauma capacity and forcing life-or-death triage decisions under extreme pressure [gunaratna2013homegrown]. Such Mass Casualty Incidents (MCIs) resulting from terrorist attacks, armed conflicts, or large-scale natural disasters can involve hundreds or thousands of victims and overwhelm local emergency services instantly [tahernejad2024application]. In these chaotic scenarios, resources are limited, communication can be disrupted, and responders are exposed to fatigue and personal risk while providing lifesaving care to multiple casualties simultaneously. The shortage of medical personnel further complicates response efforts, forcing small, exhausted teams to triage a large number of patients under hazardous conditions, where accurate patient prioritization is essential, as delays in care can be fatal. However, conventional manual triage methods can fail when faced with high casualty volumes and complex injury profiles [gabbe2022review, super1994start].
To address these challenges, recent efforts have focused on autonomous perception systems that improve situational awareness and decision support technologies. For example, robotic and drone platforms can be developed, capable of surveying disaster zones to identify and locate victims while remotely assessing their injuries and vital signs [agarwal2014characteristics]. These autonomous systems are equipped with multimodal sensors, such as radar, lidar, multispectral, thermal, and visible light cameras, that can noninvasively assess heart rate, respiratory rate, bleeding, amputations, responsiveness, and wounds. By providing live updates of these data to healthcare workers, these technologies can accelerate diagnosis and interventions.
Although these autonomous perception capabilities represent significant technological advances, the algorithms that power them face fundamental limitations that compromise their reliability in high-stakes MCI scenarios. Current models suffer from three critical limitations: algorithmic isolation-treating each vital sign independently, context blindness-ignoring clinical interdependencies, and brittleness under the noisy, partial data conditions prevalent to disaster scenarios. For instance, an elevated heart rate algorithm may indicate urgency, but without considering visible trauma or level of consciousness, it cannot distinguish between pain response, blood loss, or anxiety, and its distinction can determine survival. Robust decision-making requires fusing multimodal outputs and embedding medical expertise into the reasoning process [wadden2022defining].
@@ 2.1 Bayesian networks
Bayesian Networks (BNs) [pearl1988probabilistic], introduced in the 1980s, are Probabilistic Graphical Models (PGMs) that represent uncertain knowledge using Directed Acyclic Graphs (DAGs), where nodes denote random variables and edges signify conditional dependencies. The power of BNs stems from their ability to compactly represent a joint probability distribution over a set of variables $\mathcal{X}=\{X_{1},\dots,X_{n}\}$. The structure of the DAG encodes conditional independence assumptions, allowing for a computationally tractable factorization of the joint distribution:
@@ 2.2 Bayesian networks in medical applications
Bayesian Networks (BNs) have been effectively used for clinical triage by modeling relationships between patient symptoms and outcomes. For instance, BN-based systems have been developed for triaging non-traumatic abdominal pain [sadeghi2006bayesian] and detecting pediatric asthma with high accuracy [sanders2006prospective]. While other studies also explore BNs for triage support, they often highlight the complexity of this approach [abad2008evolution, butcher2020extending, olszewski2003bayesian].
Concurrently, recent advances in machine learning have enabled remote, non-contact prediction of vital signs, such as heart rate from eye movements [zheng2022heart], SpO2 levels [al2021non], and blood pressure via remote photoplethysmography [schrumpf2021assessment]. A recent systematic review and meta-analysis confirms the potential of this approach, showing that data-driven AI models generally outperform conventional triage tools in predicting trauma outcomes [adebayo2023exploring]. However, a significant challenge remains to translate the success of these individual models into a robust integrated system capable of assessing each casualty case as a whole. Relying solely on the isolated predictions of these diverse, often "black-box" neural networks can lead to inaccuracies, particularly in complex environments.
@@ 3.1 Environment
Table 1: Scoring criteria for emergency scenarios. Points are awarded based on the official DARPA Triage Challenge scoring rubric [dtc_resources]. Higher scores were given for correctly assessing critical conditions, such as respiratory distress and hemorrhage, within the golden window (GW), defined as the first half of the simulation scenario, due to their urgency and clinical importance. Maximum achievable score for each casualty is equal to 12 points.
@@ 3.2 System architecture and integration
To address the multifaceted issues involved in the assessment of casualties, our solution leverages a probabilistic modeling framework grounded in domain expertise. The underlying BN model was constructed and processed using GeNIe Modeler [genie2022]. Its seamless integration into the Robot Operating System 2 (ROS2) stack was achieved through a custom interface developed with SMILE [smile2022], which facilitates real-time inference and data exchange. This implementation directly leverages the principles outlined in [druzdzel1999smile] for practical application within a modern robotic middleware ecosystem.

### v2

@@ 1 Introduction
When a terrorist attack struck the Boston Marathon in 2013, emergency responders faced around 250 casualties within minutes, overwhelming the city’s trauma capacity and forcing life-or-death triage decisions under extreme pressure [11]. Such Mass Casualty Incidents (MCIs) resulting from terrorist attacks, armed conflicts, or large-scale natural disasters can involve hundreds or thousands of victims and overwhelm local emergency services instantly [18]. In these chaotic scenarios, resources are limited, communication can be disrupted, and responders are exposed to fatigue and personal risk while providing lifesaving care to multiple casualties simultaneously. The shortage of medical personnel further complicates response efforts, forcing small, exhausted teams to triage a large number of patients under hazardous conditions, where accurate patient prioritization is essential, as delays in care can be fatal. However, conventional manual triage methods can fail when faced with high casualty volumes and complex injury profiles [10, 17].
To address these challenges, recent efforts have focused on autonomous perception systems that improve situational awareness and decision support technologies. For example, robotic and drone platforms can be developed, capable of surveying disaster zones to identify and locate victims while remotely assessing their injuries and vital signs [3]. These autonomous systems are equipped with multimodal sensors, such as radar, lidar, multispectral, thermal, and visible light cameras, that can noninvasively assess heart rate, respiratory rate, bleeding, amputations, responsiveness, and wounds. By providing live updates of these data to healthcare workers, these technologies can accelerate diagnosis and interventions.
Although these autonomous perception capabilities represent significant technological advances, the algorithms that power them face fundamental limitations that compromise their reliability in high-stakes MCI scenarios. Current models suffer from three critical limitations: algorithmic isolation-treating each vital sign independently, context blindness-ignoring clinical interdependencies, and brittleness under the noisy, partial data conditions prevalent to disaster scenarios. For instance, an elevated heart rate algorithm may indicate urgency, but without considering visible trauma or level of consciousness, it cannot distinguish between pain response, blood loss, or anxiety, and its distinction can determine survival. Robust decision-making requires fusing multimodal outputs and embedding medical expertise into the reasoning process [19].
@@ 2.1 Bayesian networks
Bayesian Networks (BNs) [13], introduced in the 1980s, are Probabilistic Graphical Models (PGMs) that represent uncertain knowledge using Directed Acyclic Graphs (DAGs), where nodes denote random variables and edges signify conditional dependencies. The power of BNs stems from their ability to compactly represent a joint probability distribution over a set of variables $\mathcal{X}=\{X_{1},\dots,X_{n}\}$. The structure of the DAG encodes conditional independence assumptions, allowing for a computationally tractable factorization of the joint distribution:
@@ 2.2 Bayesian networks in medical applications
Bayesian Networks (BNs) have been effectively used for clinical triage by modeling relationships between patient symptoms and outcomes. For instance, BN-based systems have been developed for triaging non-traumatic abdominal pain [14] and detecting pediatric asthma with high accuracy [15]. While other studies also explore BNs for triage support, they often highlight the complexity of this approach [1, 7, 12].
Concurrently, recent advances in machine learning have enabled remote, non-contact prediction of vital signs, such as heart rate from eye movements [20], SpO2 levels [4], and blood pressure via remote photoplethysmography [16]. A recent systematic review and meta-analysis confirms the potential of this approach, showing that data-driven AI models generally outperform conventional triage tools in predicting trauma outcomes [2]. However, a significant challenge remains to translate the success of these individual models into a robust integrated system capable of assessing each casualty case as a whole. Relying solely on the isolated predictions of these diverse, often "black-box" neural networks can lead to inaccuracies, particularly in complex environments.
@@ 3.1 Environment
Table 1: Scoring criteria for emergency scenarios. Points are awarded based on the official DARPA Triage Challenge scoring rubric [8]. Higher scores were given for correctly assessing critical conditions, such as respiratory distress and hemorrhage, within the golden window (GW), defined as the first half of the simulation scenario, due to their urgency and clinical importance. Maximum achievable score for each casualty is equal to 12 points.
@@ 3.2 System architecture and integration
To address the multifaceted issues involved in the assessment of casualties, our solution leverages a probabilistic modeling framework grounded in domain expertise. The underlying BN model was constructed and processed using GeNIe Modeler [5]. Its seamless integration into the Robot Operating System 2 (ROS2) stack was achieved through a custom interface developed with SMILE [6], which facilitates real-time inference and data exchange. This implementation directly leverages the principles outlined in [9] for practical application within a modern robotic middleware ecosystem.